OpenStudy (anonymous):

Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (hba):

Um, Eq(2)=) 3x + y = 3 y=3-3x---(3) Now put y or Eq(3) in Eq(1) 4x + 3y = -1

terenzreignz (terenzreignz):

Hang on, @mikala1 Maybe you ought to do the 'deciding' part first?

OpenStudy (hba):

She is gonna do substitution only because she never gets the elimination @terenzreignz

terenzreignz (terenzreignz):

Well, whatever works, I guess :)

OpenStudy (anonymous):

can some one show me how to do this one in elimination

OpenStudy (hba):

@mikala1 See what i have showed.

OpenStudy (anonymous):

yes i do thank you

terenzreignz (terenzreignz):

That was substitution, though.

OpenStudy (anonymous):

could you show the elimination to me plz

terenzreignz (terenzreignz):

First, let's do a quick review of the properties of equality, just so you know what you're doing... First let's assume a = b c = d and n is some real number.

terenzreignz (terenzreignz):

First, we could multiply n to both sides of the equation, and the equality holds: a = b na = nb

terenzreignz (terenzreignz):

Also, we could add the left-hand sides of the equation, and this would be equal to the sum of the right-hand sides of the equation: a = b c = d a + c = b + d

terenzreignz (terenzreignz):

You understand these?

OpenStudy (anonymous):

yes so far

terenzreignz (terenzreignz):

Okay, we are going to apply them... Back to the system of equations... 4x + 3y = -1 3x + y = 3 The trick here is to make use of that second property I showed you, where you "add" both equations. Getting it so far?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

We *could* go ahead and add both equations outright : (4x + 3y) + (3x + y) = -1 + 3 7x + 4y = 2 But this doesn't really help us, right?

OpenStudy (anonymous):

no not really

terenzreignz (terenzreignz):

This is where the first property comes in :) We want to make it so that adding the equations gets rid of one of the unknowns. Let's rewrite that system: 4x + 3y = -1 3x + y = 3 We want to get rid of one of these unknowns (either x or y) In time, you'll learn to recognise which is easier to eliminate, but for now, take my word for it that y is easier to eliminate. Getting it so far?

OpenStudy (anonymous):

yes thank you keep going i will say if i dont understand

terenzreignz (terenzreignz):

I'd prefer you solve this yourself, though. I'll guide you through it. Again, rewriting the system: 4x + 3y = -1 3x + y = 2 We want to get rid of y, right? Look at the first equation. What do you add to 3y so that you get zero?

OpenStudy (anonymous):

you will add -1 right?

terenzreignz (terenzreignz):

No. My question was What was it you add to 3y so that you get 0... In symbols 3y + ? = 0 Is it clearer now?

OpenStudy (anonymous):

im not sure what you add i thought the teacher siad -1 or 4 x

terenzreignz (terenzreignz):

well, you could just sort of find the value of that "?" there 3y + ? = 0 and you bring 3y to the other side to get ? = -3y Understand it now? :)

OpenStudy (anonymous):

yes so its 3

terenzreignz (terenzreignz):

not just 3 you add -3y or in other words you subtract 3y Get it?

OpenStudy (anonymous):

yes i for got the y sorry about that

terenzreignz (terenzreignz):

Back to the system: 4x + 3y = -1 3x + y = 2 You know that in the first equation, you need to somehow add "-3y" to get rid of the y, right?

OpenStudy (anonymous):

yes and i only needed the fisrt one helped on

terenzreignz (terenzreignz):

So, that's your hint. You have to somehow acquire a "-3y" at your second equation. How do you go about that? (HINT: Use the first property of equality we discussed)

OpenStudy (anonymous):

ok thank you so much

terenzreignz (terenzreignz):

You already get it? Brilliant :)

OpenStudy (anonymous):

yes i did thank you